OpenStudy (anonymous):

What is the degree measure of an arc 4 ft. long in a circle of radius 10 ft.?

OpenStudy (jdoe0001):

\(\bf \textit{arc's length}=s=\cfrac{r{\color{blue}{ \theta}}\pi}{180}\qquad \begin{cases} s=4\\ r=10 \end{cases}\qquad 4=\cfrac{10{\color{blue}{ \theta}}\pi}{180}\) solve for the angle, that is \(\color{blue}{ \theta}\)